Problem 4. Two hockey teams, A and B play a series of games, until one of the teams wins 4 games. Suppose team A has probability p of winning each game, and games are independent. Let X be the total number of games that are played (a) Find the probability mass function of X (b) What is the probability that team A wins the series conditioned on X-47 (c) What is the probability that team A wins the series conditioned on X 7? (Simplify your expressions as much as possible.)
Answer #1

a)pmf of X is as follows:

P(X=4)=P(A wins first 4 games)+P(B wins 4 first games)=p4+(1-p)4

P(X=5)=P(A wins 3 of first 4 and wins 5th game)+P(B wins 3 of first 4 and wins 5th game)

=4C3p4(1-p)1+4C3p1(1-p)4 =4*p*(1-p)*(p3+(1-p)3)

P(X=6)=5C3p4(1-p)2+5C3p2(1-p)4 =10*p2(1-p)2*(p2+(1-p)2)

P(X=7)=6C3p4(1-p)3+6C3p3(1-p)4 =20*p3(1-p)3*(p+(1-p))=20*p3(1-p)3

b)

P(Awins |X=4)=P(A wins first 4 games)/P(X=4)=p4/(p4+(1-p)4)

c)

P(Awins |X=7)=6C3p4(1-p)3/(20*p3(1-p)3) =p